# Are Image Distributions Indistinguishable to Humans Indistinguishable to Classifiers?

### v1

@@ 5.2 Self-supervised classifiers can identify the generated images
Despite the remarkable performance of diffusion models in generating realistic images, classifiers easily distinguish real from generated images with far fewer training samples than required for the diffusion models, achieving high accuracy across various combinations. Besides, self-supervised classifiers also demonstrate the ability to identify generated images.
@@ 6 Classifier contradictions with FID and human judgments
Previously, we demonstrated that classifiers can effortlessly distinguish between real and generated images. In this section, we explore the intriguing contradictions between classifiers and both FID and human judgments through two sets of experiments. The first experiment involves EDM2-XS and EDM2-XXL, which share identical network architectures and training methodologies but differ in the number of parameters, resulting in FID scores of 2.91 and 1.81, respectively. The second experiment includes DiT-XL, U-ViT-H/2, and EDM2-XXL, which, despite having similar FID scores, differ in network architecture, sampling methods, and training protocols.
Table 4: Distribution classification accuracy on combinations of different diffusion models with similar FIDs. An intriguing phenomenon occurs when more distributions are added: the classifier achieves higher distribution classification accuracy.
Combinations Model Accuracy
@@
Figure 3: User study. Classifiers can distinguish between different diffusion models with similar FIDs but struggle with variations of the same diffusion model differing in parameters. In contrast, humans show the opposite tendency.
6.1 User study
Here, we conduct a user study to explore human capabilities in distinguishing generated distributions from real ones and classifying from which diffusion model an image is generated (see Fig. 10). Details of the study can be found in Appendix C.4. The results are presented in Fig. 3, with classifier performance represented by the median accuracy achieved using three architectures. Specifically, humans are unable to distinguish between different diffusion models with similar FIDs, yet they demonstrate some capability in distinguishing different variants of EDM2.
6.2 Classifiers distinguish between different diffusion models with similar FIDs
@@
As illustrated in Tab. 4, we present the distribution classification accuracy for three combinations of datasets generated by diffusion models with similar FIDs. We observe that classifiers can effortlessly distinguish between different diffusion models, even those with closely matched FID scores, while humans cannot. Notably, an intriguing phenomenon occurs when more datasets are added: the classifier achieves even higher classification accuracy. This phenomenon differs from the findings in  [28]. A possible explanation is that all these datasets are related to ImageNet, whether directly from ImageNet or from diffusion models trained on it. Thus, adding new distributions enables the classifier to apply newly learned patterns to enhance the classification accuracy of previous combinations.
6.3 Classifiers struggle to distinguish different models in the same family
As shown in Tab. 5, we present the distribution classification accuracy for combinations of EDM2-XS and EDM2-XXL, which are the smallest and largest models in the EDM2 family, respectively. We observe that classifiers struggle to distinguish between different models within the same family. Only the ResNet-50 demonstrates some ability to classify them, while the other two models perform close to random selection. Notably, humans also show some ability to classify them (see Fig. 3).
Conclusion 2.
Classifiers conflict with FID and human assessments in distribution classification. They can distinguish between different diffusion models with similar FID but different inductive biases but struggle to do so within the same family, where models share inductive biases but differ in visual quality. In contrast, humans show a different trend.
7 How classifiers distinguish between different distributions
The previous findings illustrate the classifier’s capability to differentiate between real and generated images and detect different inductive biases in diffusion models. In this section, we delve into how classifiers distinguish between different distributions to explore the reasons behind this capability.
7.1 Classifier uses edge information to distinguish real from generated images
We first use Grad-CAM [31] to investigate which aspects of image classifiers depend on to distinguish between real and generated images. We use the combination of U-ViT-H/2 and ImageNet-256, with ResNet-50 as the classifier architecture. We observe that classifiers rely mainly on edge information of an image to make their classification, regardless of whether the image is real or generated (see Fig. 4, first column). This led us to question whether the classifier’s ability to distinguish between real and generated images would persist if edge information were reduced. To validate this, we center-crop images from a resolution of 256 to smaller sizes of 128, 64, 32, and 16, respectively, and train the classifier from scratch. The results, shown in Tab. 6, indicate that even without original edge information, the classifier can also distinguish between real and generated images with high accuracy based on the new edge information presented at each lower resolution (see Fig. 4).
@@ 7.1 Classifier uses edge information to distinguish real from generated images
Table 5: Classifiers struggle to differentiate between the smallest and largest models of EDM2. Despite notable differences in FID, the classifiers show limited ability to accurately distinguish between different EDM2 variants.
@@
Figure 4: Grad-CAM [31] results on validation images (top) and samples of U-ViT-H/2 (bottom).
7.2 Classifiers can discern real from generated images using any part
Building on our findings that classifiers can effectively distinguish between real and generated images using just the central portion, we extended our investigation to see if classifiers can make distinctions using any arbitrary part of an image. To explore this, we focused on two resolutions, 128 and 64, and for each resolution, we randomly cropped images four times to train the classifiers (see Fig. 7 in the Appendix for visualization of crops). For images cropped to the resolution of 128 and 64, classifiers achieve an accuracy of $99.97\pm 0.004$ and $99.97\pm 0.007$, respectively (see Tab. 6). The quantitative results on Tab. 6 demonstrate that, to a certain extent, classifiers are capable of classifying whether an image is real or generated based on any part of the image.
@@ 7.2 Classifiers can discern real from generated images using any part
Figure 5: Classifier achieves high accuracy across various combinations with limited frequency components, except when only a little low-frequency components is present.
7.3 The impact of frequency on distribution classification task
Having established that classifiers mainly rely on edge information to classify both original and cropped images, regardless of the position and size of the cropped images, we will next explore the impact of frequency on classification. To this end, we preprocess the images with low-pass, high-pass and band-pass filters as defined in Sec. 2.2. For low-pass and high-pass filters, we apply four thresholds: 10, 30, 50 and 100. For the band-pass filter, we choose four threshold intervals: 0-10, 10-30, 30-50, and 50-100. We visualize an example of original and processed images in Fig. 6.
@@ 7.3 The impact of frequency on distribution classification task
Figure 6: Visualization of frequency domain processing. Top: Low-Pass, Middle: High-Pass, each processed with increasing thresholds: 10, 30, 50, 100. Bottom: Band-Pass, processed with band thresholds: 0-10, 10-30, 30-50, 50-100.
As shown in Fig. 5, using ResNet-50 as the classifier architecture, we observe two intriguing phenomena when classifying images as either generated or real, based on different frequency filters. First, classifiers can effortlessly classify images even when primarily high-frequency components are present (i.e., threshold at 100). Second, classifiers can classify images even with any interval of frequency components, except when only very little low-frequency information is available.
Moreover, we also conduct an experiment on the combinations of EDM2-XS and EDM2-XXL (see details in Appendix. C.3), we observe that for models within the same diffusion model family, which share inductive biases but differ in visual quality, classifiers unable to distinguish between them based on any specific frequency band, achieving accuracy close to random guessing (around 50%).
Conclusion 3.
Classifiers primarily rely on edge information to distinguish between real and generated images. Even when original edges are removed via center cropping, they utilize new edges at lower resolutions. Additionally, classifiers can distinguish based on any part and maintain high accuracy unless only minimal low-frequency information is available. In contrast, for models within the same diffusion family, classifiers’ accuracy drops close to random guessing (around 50%) when distinguishing based on any specific frequency bands.
8 Conclusions
Our paper proposes the distribution classification tasks to explore a fundamental question: How far is the generated distribution from the real distribution? Our study reveals that classifiers can effectively distinguish between diffusion-generated and real distributions, despite the low FID scores and lifelike appearance of the generated images. Moreover, classifiers are more sensitive to the inductive biases of different diffusion models than humans, who rely more on visual quality. Finally, classifiers primarily use edge information for these distinctions, maintaining high accuracy even with limited frequency components, except when only minimal low-frequency components are present.
Limitations and social impact. Our findings are based on specific datasets and model architectures, which may limit their generalizability. Future work can explore a wider range of datasets to validate our results. Further, our paper may draw attention to enhancing the edge or high-frequency details in generated images in the future, potentially making "DeepFakes" more difficult to detect.
@@ Appendix A Detail experiment settings
We implement our experiments upon the official code of ConvNeXt [53], U-ViT [20], DiT [21], EDM [26], EDM2 [19]. We also utilize the official models of Pixart-$\alpha$ [23], SDXL [24], and Playground-v2.5 [22]. The respective links and licenses are detailed in Tab. 7.
@@
Dataset. For label-to-image, we consider the CIFAR [48] and ImageNet [49] datasets, which are well-established and widely recognized benchmarks in the field of image generation. For text-to-image, we utilize the COCO2014 dataset [50], known for its rich annotations and diverse image content. For CIFAR-10, we use the real CIFAR-10 training set and validation set to construct the real distribution and use the diffusion model to generate 50k training images and 10k validation images to construct the generated distribution. In the case of ImageNet, we consider 256 and 512 resolutions, which are common resolutions for ImageNet image generation tasks. We randomly sample 100k training images and 50k validation images from the training set and the validation set of ImageNet to construct the real distribution and use the diffusion models to generate 100k training images and 50k validation images to construct the generated distribution. For the real dataset, we adopt the data processing method from ADM [51] to modify the ImageNet dataset into two common resolutions: ImageNet-256 and ImageNet-512, where the numbers indicate the resolution of the data. In addition, for the COCO2014 dataset, by default, we construct the real distribution by randomly sampling 10k training images and 1k validation images from the respective training and validation sets of COCO2014. Each image in this dataset is associated with five captions. To create the generated distribution, we randomly select one of the five captions for each real image to serve as a prompt. These prompts are then used by the diffusion models to generate an equivalent number of images to construct the generated distribution.
@@
Classifier. By default, we employ the ResNet-50 [52] as the classifier architecture. For completeness, we also consider ConvNeXt-T [53] and ViT-S [54]. Our pre-processing protocol follows the standard supervised training approach [53]. Specifically, during training, classifiers process randomly augmented crops of $224\times 224$ images. During validation, images are resized so that their smaller dimension reaches 256 pixels while preserving the original aspect ratio. Subsequently, these images are center cropped to $224\times 224$ pixels before being fed into the model. For the experiments on CIFAR-10, we initialize our classifier with the ResNet-50, pre-trained on ImageNet. For completeness, we also present the result trained from scratch (see experiments in Appendix. C.1). In the case of ImageNet, we opt to train the ResNet-50 model from scratch.
@@
Diffusion model. For the generation of CIFAR-10, we consider two diffusion models: EDM [26] and U-ViT [20]. Both models have demonstrated strong generation performance on the CIFAR-10 [48]. Quantitatively, EDM achieves an FID of 1.79, and U-ViT achieves an FID of 3.11. To improve efficiency, we modified U-ViT’s sampling method from Euler-Maruyama to DPM-Solver [57] and reduced the sampling steps from 1,000 to 50. These adjustments resulted in U-ViT achieving an FID of 3.65 on CIFAR-10. For the generation of ImageNet-256, we explore three diffusion models: EDM2 [19], U-ViT-H/2 [20], DiT-XL [21]. Both DiT and U-ViT are prominent diffusion transformer architectures known for their scalability and strong performance. U-ViT-H/2 achieves an FID of 2.29 on ImageNet-256, and DiT-XL/2 achieves an FID of 2.27. We consider EDM2 to incorporate a UNet-based architecture, which was traditionally used before the rise of diffusion transformers. Since EDM2 is originally designed for ImageNet-512 generation, we resize the generated images from 512 to 256 resolution to suit our ImageNet-256 task. In this way, EDM2-XXL achieves an FID of 2.14 on this task, which is similar to the FID achieved by U-ViT and DiT. For the generation of ImageNet-512, we use EDM2 [19], which achieves state-of-the-art performance on this task with an FID of 1.81. For the generation of COCO, we consider three state-of-the-art text-to-image diffusion models: Pixart-$\alpha$ [23], SDXL [24], and Playground-v2.5 [22].

### v2

@@ 5.3 Self-supervised classifiers can also identify generated images
Despite the success of diffusion models in image generation, classifiers easily distinguish real from generated distribution, achieving high accuracy across various combinations. Besides, self-supervised classifiers also effectively identify generated distribution.
@@ 6 Classifier contradictions with FID and human judgments
Previously, we demonstrated that in most settings, classifiers can effortlessly distinguish between real and generated distributions, indicating that the generated distribution is still far from the real one. In this section, we explore the intriguing contradictions between classifiers and widely used evaluation protocols, such as FID and human judgments, through two sets of experiments. The first experiment compares EDM2-XS and EDM2-XXL, which share the same network architectures and training methodologies but differ in the number of parameters, resulting in FID scores of 2.91 and 1.81, respectively. The second experiment examines DiT-XL, U-ViT-H/2, and EDM2-XXL, which, despite having similar FIDs, vary in network architecture, sampling methods, and training protocols.
Table 5: Distribution classification accuracy for various combinations of diffusion models with similar FIDs and ImageNet. Notably, classifier accuracy improves as more distributions are added. Here, combinations refer to multiple distributions, where the number of distributions corresponds to the number of classes in distribution classification.
Distribution combinations Model Accuracy
@@
Figure 3: User study. Classifiers can easily distinguish between diffusion models with similar FIDs (U-H/2, D) but struggle with models from the same family that differ in parameters (E2-XS, E2-XXL). In contrast, humans show the opposite trend.
To analyze human judgments, we conducted a user study to evaluate the human ability to distinguish between different distributions (see the user interface in Fig. 10). Further details of the study are provided in Appendix C.6. The results are shown in Fig. 3, where individual variability is represented using error bars, indicating the standard deviation across participants/classifier performance.
In our first set of experiments, we analyze models within the same family: EDM2-XS and EDM2-XXL, the smallest and largest models in the EDM2 family, respectively. As shown in Tab. 6, we report the distribution classification accuracy for distinguishing between EDM2-XS and EDM2-XXL. The results show that classifiers struggle to differentiate between models from the same family. Only ResNet-50 shows some ability to distinguish them, while the other classifiers perform close to random chance. This contrasts sharply with the task of distinguishing between ImageNet-512 and EDM2-XXL (in Tab. 2), where all classifiers achieve over 86% accuracy. Thus, compared to distinguishing between diffusion models and the real distribution, neural networks view models within the same family—due to their shared inductive biases—as learning similar distributions, making them difficult to distinguish. Furthermore, in our user study (see Fig. 3), unlike the other tasks where classifiers significantly outperform humans, in the task of distinguishing between models within the same family, humans perform comparably to the classifiers. This observation highlights a limitation of neural network-based classifiers: despite the noticeable difference in FID scores between EDM2-XS and EDM2-XXL (2.91 and 1.81, respectively), the classifiers struggle to differentiate between them.
@@
In our second group of experiments, we analyze different diffusion models with similar FIDs. As shown in Tab. 5, we present the distribution classification accuracy for three combinations of diffusion models with similar FIDs. We observe that classifiers can easily distinguish between these models, achieving extremely high accuracy, while humans perform close to random guessing (see Fig. 3). This highlights a key advantage of classifiers over widely used evaluation protocols: despite the FID scores being very close and humans being unable to differentiate between the models, classifiers can easily identify the differences. Moreover, an intriguing phenomenon occurs when additional distributions are introduced: the classifier’s accuracy improves further. This contrasts with the findings in Liu & He (2024). A possible explanation is that all of these distributions are either directly related to ImageNet or generated by diffusion models trained on it. As new distributions are added, the classifier can leverage newly learned patterns to enhance the classification accuracy for previous combinations.
@@
Table 6: Classifiers struggle to differentiate between the smallest and largest EDM2 models. Despite notable differences in FID (E2-XS: 2.91, E2-XXL:1.81), the classifiers show limited ability to accurately distinguish between these EDM2 variants.
@@
Conclusion 2.
Classifiers conflict with FID and human assessments in distribution classification. They can distinguish between different diffusion models with similar FID but struggle to do so within the same family. In contrast, humans show a different trend.
7 Implications of the distribution classification task
Previously, we demonstrated the classifier’s ability to distinguish real from generated distributions, as well as the contradictions between classifiers and evaluation methods like FID and human judgments. In this section, we show some important implications of the distribution classification task.
7.1 Diagnosing the problem within diffusion models
A natural implication of classifiers is to diagnose the problem within diffusion models. Our approach complements traditional metrics like FID and human judgments by offering a more challenging evaluation method. It can also be extended to analyze specific features of generated data, providing deeper insights into the intrinsic characteristics of diffusion models, such as spatial and frequency features. In this section, we focus on advanced diffusion transformers, specifically U-ViT-H/2 (Bao et al., 2023), and investigate two key aspects: spatial features and frequency ranges. Our objective is to identify which spatial features diffusion models learn well, making them harder to distinguish, and which they struggle with, making them easier to detect. Additionally, we assess performance across different frequency ranges to pinpoint where diffusion models succeed and where they fail. For this analysis, we use U-ViT-H/2 trained on ImageNet-256, with ResNet-50 as the classifier.
@@ 7.1 Diagnosing the problem within diffusion models
Figure 4: Classifier achieves high accuracy across various combinations with limited frequency components, except when only minimal low-frequency components are present (in subfig (a) low-pass filter threshold = 10, in subfig (c) band-pass filter threshold: 0-10).
@@
Figure 5: Visualization of frequency domain processing. Top: Low-pass filters, Middle: High-pass filters, each processed with increasing thresholds: 10, 30, 50, 100. Bottom: Band-pass filters, processed with band thresholds: 0-10, 10-30, 30-50, 50-100.
In our first set of experiments, we focus on the spatial features. We start by progressively reducing central information through center-cropping images from a resolution of 256 down to 128, 64, 32, and 16, followed by training the classifier from scratch. The results, shown in Tab. 7, indicate that even with minimal central information, the classifier can accurately distinguish between real and generated distributions. We then assess whether classifiers can differentiate based on any arbitrary part of the image. We cropped images at resolutions of 128 and 64, and randomly selected four sections per image for training (see Fig. 7 for visualizations). The classifiers achieved accuracies of $99.97\pm 0.004$% for 128 and $99.97\pm 0.007$% for 64 (see Tab. 7). These results demonstrate that even with the advanced diffusion transformer U-ViT-H/2, which achieves an FID of 2.29, classifiers can reliably distinguish between real and generated images using any spatial part of the image.
In our second set of experiments, we focus on the frequency aspect. We preprocess images using low-pass, high-pass, and band-pass filters as described in Sec. 2.2. For the low-pass and high-pass filters, we apply thresholds of 10, 30, 50, and 100, while for the band-pass filter, we use frequency intervals of 0-10, 10-30, 30-50, and 50-100. Fig. 5 provides examples of the original and filtered images. As shown in Fig. 4, using ResNet-50 as the classifier, we observe that classifiers struggle to maintain high accuracy when low-frequency information is minimal but perform well across all other frequency intervals, even with only high-frequency components (e.g., threshold of 100). These results suggest that diffusion models effectively learn low-frequency features, making it more difficult for classifiers to distinguish real from generated images in this range. However, in higher frequency bands, diffusion models perform worse, making it easier for classifiers to identify generated images.
7.2 Evaluating the use of generated data
Another natural implication is using classifiers to evaluate the use of generated data. Our previous findings show a significant difference between generated and real distributions, raising the question of whether this mismatch causes undesirable behaviors in models trained on synthetic data. One well-known phenomenon is that training generative models on data produced by previous generations in a continuous loop leads to a gradual decline in output quality and diversity (Shumailov et al., 2024; Alemohammad et al., 2024), ultimately resulting in model collapse, known as Model Autophagy Disorder (MAD). This is surprising, as data generated by modern diffusion models often exhibit high realism. However, using this data to train next-generation models triggers MAD. Our approach provides insight into this issue: although these models generate realistic images, the distributions they produce remain distinct from the real distribution—especially when viewed by classifiers—leading to a distribution mismatch. As this mismatch accumulates, the generated distribution drifts further from the original. To address this, Alemohammad et al. (2024) propose incorporating fresh real samples into the training data for each generation, rather than relying solely on generated data.
To better understand this phenomenon, we further explore the impact of using generative model data in downstream tasks. Specifically, we investigate whether the distribution mismatch between real and generated data causes undesirable behaviors when replacing real data with generated data in tasks such as classification. To analyze this, we use EDM (Karras et al., 2022) to generate the same number of samples as CIFAR-10, creating a synthetic dataset, EDM_S, and perform both supervised and semi-supervised classification to assess the impact. For supervised classification, we train the classifier with ConvNeXt settings (Liu et al., 2022) for CIFAR-10, using the original test set for evaluation. To create the training dataset, we mix EDM_S and CIFAR-10 with a ratio $\alpha$. We employ two strategies: one replaces $\alpha$ of CIFAR-10 with EDM_S, where $\alpha$ represents the replacement ratio, and the other augments EDM_S to CIFAR-10. As shown in Fig. 6, we observe that due to the distribution mismatch, simply replacing samples leads to a decline in accuracy. As $\alpha$ increases from 0 to 1, the top-1 accuracy gradually declines. In contrast, when augmenting EDM_S to CIFAR-10, accuracy generally improves for most $\alpha$ values. For semi-supervised classification, we use FreeMatch (Wang et al., 2023a) and FlexMatch (Zhang et al., 2021). As shown in Tab. 8, simply replacing CIFAR-10 with EDM_S reduces performance: FlexMatch’s error rate increases from 4.97$\pm$0.06to 5.85$\pm$0.02%, and FreeMatch’s from 4.90$\pm$0.04to 5.94$\pm$0.10%, consistent with the results in supervised classification. Meanwhile, DPT (You et al., 2024) augments CIFAR-10 with EDM-generated data, training with FreeMatch, and achieves a lower error rate of 4.68$\pm$0.17%.
From the analysis above, we observe that the distribution mismatch between generated and real data leads to undesirable behaviors, such as "MAD" or performance degradation, when real data is simply replaced with generated data. However, augmenting real data with generated data is more effective.
Figure 6: Top-1 accuracy on the CIFAR-10 test set: Replacement vs. Augmentation. $\alpha$ represents the ratio of EDM_S mixed with CIFAR-10.
Table 8: Augmenting real data with generated data is more effective than replacing it. indicates results from training with CIFAR-10 (real data); and indicate results where CIFAR-10 is replaced or augmented with EDM_S, respectively.
Method Error rate $\downarrow$
given ##\## labels per class (label fraction) 4 (0.08%)
FlexMatch(Zhang et al., 2021) 4.97$\pm$0.06
FlexMatch 5.85$\pm$0.02
FreeMatch(Wang et al., 2023a) 4.90$\pm$0.04
FreeMatch 5.94$\pm$0.10
FreeMatch(You et al., 2024) 4.68$\pm$0.17
8 Conclusion
Our paper proposes the distribution classification tasks to explore a fundamental question: How far is the generated distribution from the real distribution? Our findings show that classifiers can effectively distinguish between diffusion-generated and real distributions, even when diffusion models achieve low FID scores and generate lifelike images. Additionally, our study shows contradictions between classifier performance and evaluation metrics such as FID and human judgments. Finally, our methodology can naturally serve as a diagnostic tool for diffusion models, offering a reasonable explanation for Model Autophagy Disorder (MAD) and evaluating the use of generated data.
Ethics Statement
This paper enhances the understanding of generated data and the differences between generated and real distributions, including shedding light on the model autophagy disorder phenomenon and evaluating the use of generated data. However, our findings could be used to improve the realism of generated images, potentially making "DeepFakes" more difficult to detect if misused. This issue can be mitigated by advancing automatic detection methods, which is an active area of research.
Reproducibility Statement
Our submission includes PyTorch code to ensure research reproducibility. Please refer to README.md for detailed instructions such as setting up the Anaconda environment. The datasets used are public and detailed in Appendix A. Pretrained models and diffusion sampling implementations are accessible via the code links listed in Table 9, including scripts for label-to-image generation (e.g., U-ViT, DiT, EDM) and text-to-image generation (e.g., Pixart-$\alpha$, SDXL, Playground-v2.5).
@@ Appendix A Detail experiment settings
We implement our experiments upon the official code of ConvNeXt (Liu et al., 2022), U-ViT (Bao et al., 2023), DiT (Peebles & Xie, 2023), EDM (Karras et al., 2022), EDM2 (Karras et al., 2023). We also utilize the official models of Pixart-$\alpha$ (Chen et al., 2023), SDXL (Podell et al., 2023), and Playground-v2.5 (Li et al., 2024). The respective links and licenses are detailed in Tab. 9.
@@
Dataset. For label-to-image, we consider the CIFAR (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009) datasets, which are well-established and widely recognized benchmarks in the field of image generation. For text-to-image, we utilize the COCO2014 dataset (Lin et al., 2014), known for its rich annotations and diverse image content. For CIFAR-10, we use the real CIFAR-10 training set and validation set to construct the real distribution and use the diffusion model to generate 50k training images and 10k validation images to construct the generated distribution. In the case of ImageNet, we consider 256 and 512 resolutions, which are common resolutions for ImageNet image generation tasks. We randomly sample 100k training images and 50k validation images from the training set and the validation set of ImageNet to construct the real distribution and use the diffusion models to generate 100k training images and 50k validation images to construct the generated distribution. For the real dataset, we adopt the data processing method from ADM (Dhariwal & Nichol, 2021) to modify the ImageNet dataset into two common resolutions: ImageNet-256 and ImageNet-512, where the numbers indicate the resolution of the data. In addition, for the COCO2014 dataset, by default, we construct the real distribution by randomly sampling 10k training images and 1k validation images from the respective training and validation sets of COCO2014. Each image in this dataset is associated with five captions. To create the generated distribution, we randomly select one of the five captions for each real image to serve as a prompt. These prompts are then used by the diffusion models to generate an equivalent number of images to construct the generated distribution.
@@
Classifier. By default, we employ the ResNet-50 (He et al., 2016) as the classifier architecture. For completeness, we also consider ConvNeXt-T (Liu et al., 2022) and ViT-S (Dosovitskiy et al., 2020). Our pre-processing protocol follows the standard supervised training approach (Liu et al., 2022). Specifically, during training, classifiers process randomly augmented crops of $224\times 224$ images. During validation, images are resized so that their smaller dimension reaches 256 pixels while preserving the original aspect ratio. Subsequently, these images are center cropped to $224\times 224$ pixels before being fed into the model. For the experiments on CIFAR-10, we initialize our classifier with the ResNet-50, pre-trained on ImageNet. For completeness, we also present the result trained from scratch (see experiments in Appendix. C.2). In the case of ImageNet, we opt to train the ResNet-50 model from scratch.
@@
Diffusion model. For the generation of CIFAR-10, we consider two diffusion models: EDM (Karras et al., 2022) and U-ViT (Bao et al., 2023). Both models have demonstrated strong generation performance on the CIFAR-10 (Krizhevsky et al., 2009). Quantitatively, EDM achieves an FID of 1.79, and U-ViT achieves an FID of 3.11. To improve efficiency, we modified U-ViT’s sampling method from Euler-Maruyama to DPM-Solver (Lu et al., 2022) and reduced the sampling steps from 1,000 to 50. These adjustments resulted in U-ViT achieving an FID of 3.65 on CIFAR-10. For the generation of ImageNet-256, we explore three diffusion models: EDM2 (Karras et al., 2023), U-ViT-H/2 (Bao et al., 2023), DiT-XL (Peebles & Xie, 2023). Both DiT and U-ViT are prominent diffusion transformer architectures known for their scalability and strong performance. U-ViT-H/2 achieves an FID of 2.29 on ImageNet-256, and DiT-XL/2 achieves an FID of 2.27. We consider EDM2 to incorporate a UNet-based architecture, which was traditionally used before the rise of diffusion transformers. Since EDM2 is originally designed for ImageNet-512 generation, we resize the generated images from 512 to 256 resolution to suit our ImageNet-256 task. In this way, EDM2-XXL achieves an FID of 2.14 on this task, which is similar to the FID achieved by U-ViT and DiT. For the generation of ImageNet-512, we use EDM2 (Karras et al., 2023), which achieves state-of-the-art performance on this task with an FID of 1.81. For the generation of COCO, we consider three state-of-the-art text-to-image diffusion models: Pixart-$\alpha$ (Chen et al., 2023), SDXL (Podell et al., 2023), and Playground-v2.5 (Li et al., 2024).